OpenStudy (anonymous):

Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive?

OpenStudy (anonymous):

Total number of votes = 125000 So number of votes Sally received = 0.25 * 125000 = 31250 So you actually don't need the equation for this. Number of votes Maya received = 125000 - 31250 = 93750 votes.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome.